OpenStudy (anonymous):

The distance between Miami and Oriando is about 220 mile .A pilot flying from Miami to Oriando start the flight 10 degree off course to avoid a stom. a) After flying in this direction for 100 mile , how far is the plane from Orlando? b)If he pilot adjust his course after 100 miles , how much father is the flight than a direct route?

OpenStudy (anonymous):

 100/1=x/0.985, 98.5=x. So the distance from Orlando is 220-98.5=121.5mi